If in a stationary wave the amplitude at antinode is 4cm. The amplitude at midway of node and antinode is?
 

Grade:

2 Answers

Arun
5 years ago
A node has a zero amplitude and an anti-node has a maximum amplitude. So, if the amplitude at anti-node is 4 cm, therefore the exact midway between node and anti-node will be at 2 cm.
    yukti arya
    13 Points
    5 years ago
    A/=2ASin kx
    k= 2 pie/lambda
    x=lambda/8    [distance b/w node and antinode = lamda/2  … so distance b/w amplitude and 0 =lambda/4.....and mid point of these two points= lambda/8]
    2A=4 [given]
    4Sin [2pie/lambda.×lambda/8]4sin[pie/4]= 4/21/2  = 2*21/2cm